OpenStudy (lgbasallote):

\(\int (\tan x)^4 (\sec x)^2 xdx\) can you show me how it's done @amistre64 @Zarkon ? i'd love to learn that lol

OpenStudy (anonymous):

missed the other x outside, damn

OpenStudy (lgbasallote):

is it something like \(\int u^4 \tan^{-1} u du\)?

OpenStudy (anonymous):

Yes, now you can use by parts.

OpenStudy (lgbasallote):

haha we were doing this in another problem..it was just a typo of an xdx there but amistre and zarkon wanted to do it so i typed it in :P

OpenStudy (amistre64):

i was gonna do it by parts: \[\int t^4s^2xdx=x\frac1{5}t^5-\frac{1}{5}\int t^5 dx \] \[\int t^4s^2xdx=x\frac1{5}t^5-\frac{1}{5}\int t^3(s^2+1) dx \] \[\int t^4s^2xdx=x\frac1{5}t^5-\frac{1}{5}\int t^3s^2+t^3 dx \] \[\int t^4s^2xdx=x\frac1{5}t^5-\frac{1}{5}(\frac{1}{4}t^4+\int t^3 dx) \]

OpenStudy (amistre64):

\[\int t^3dx=\int t(ss+1)dx\] \[=\int (s)ts+t dx\] \[=\frac{1}{2}s^2+\int t dx\] \[=\frac{1}{2}s^2-ln(cos)+c\]

OpenStudy (lgbasallote):

huh..didnt know it was this complicated :P hahaha

OpenStudy (amistre64):

not complicated, just drawn out :)

OpenStudy (lgbasallote):

would \(\int u^4 \tan^{-1}udu\) work too?

OpenStudy (amistre64):

you could give it a shot and see :)

OpenStudy (amistre64):

down the tan-1 to a function without a trig

OpenStudy (chrisplusian):

my son calls me a math nerd for taking honors classes and being concerned that I am not at an A in the class right now. He says dad your a math nerd because a B isn't good enough. He should read this....... he would think I was normal:) keep up the good work you guys

OpenStudy (amistre64):

:) you have to take classes to learn this stuff? lol

OpenStudy (chrisplusian):

Nope!

OpenStudy (amistre64):

im saving my B for english class ;)

OpenStudy (lgbasallote):

hmm downing tan^-1 seems hard -__- i'll just take your method amistre

OpenStudy (chrisplusian):

I won't take a B in anything ..... well not yet. And I would definitly not take a B in a math

OpenStudy (chrisplusian):

class

OpenStudy (amistre64):

D tan-1 = 1/(1+u^2) u up to u^5/5 integrate u^5/5(1+u^2) long divide it and see what we get

OpenStudy (amistre64):

u^3 -u + u/(u^2+1) -------------------- u^2 + 1 ) u^5 (u^5+u^3) -u^3 (-u^3-u) u

OpenStudy (lgbasallote):

you used tan^-1 as u and u^4 as dv right?

OpenStudy (amistre64):

yes

OpenStudy (lgbasallote):

and this one is the uv part right...

OpenStudy (amistre64):

v du yes

OpenStudy (lgbasallote):

oh vdu...so combined it's \(\LARGE (\tan^{-1} u)(\frac{u^5}{5}) - [u^3 - u + \frac{u}{(u^2 + 1)}]\) right?

OpenStudy (amistre64):

yes, but finish integrateing the stuff after the - sign

OpenStudy (lgbasallote):

oh oh yeah it's integral...this is seriously complicated lol i see another u^2 + 1 so im guessing another arctan @_@ i'll just go with your method

OpenStudy (amistre64):

\[...=tan^{-1}(u)\frac{u^5}{5}-\frac{1}{5}\int u^3-u+\frac{u}{u^2+1}dx\]

OpenStudy (amistre64):

that actually lns up since we can create a du above it ....

OpenStudy (lgbasallote):

arggghhh this is a pain in the head i quit lol i'll use your method

OpenStudy (amistre64):

\[...=tan^{-1}(u)\frac{u^5}{5}-\frac{1}{5}\left(\frac{u^4}{4}-\frac{u^2}{2}+\frac{1}{2}ln(u^2+1 )\right)\]

OpenStudy (amistre64):

as long as i kept it all straight

OpenStudy (lgbasallote):

stop showing me these T_T i dont want to see anymore